OpenStudy (anonymous):

is 8d-4d-6d-8=2d no solution?

OpenStudy (saifoo.khan):

\[8d - 4d - 6d - 8 = 2d\]Simplify..

OpenStudy (saifoo.khan):

\[(8-4-6)d - 8 = 2d\] \[-2d - 8 = 2d\] Solve for d.

OpenStudy (anonymous):

No solution

OpenStudy (anonymous):

@Abarret1115 @saifoo.khan i thought it was no solution but then ... idk

OpenStudy (saifoo.khan):

I did a typo! -2d - 2d = 8 -4x = 8 And so on..

OpenStudy (saifoo.khan):

There is a solution

OpenStudy (anonymous):

-2 thats what i thought

OpenStudy (anonymous):

thank yaaaaaa